OpenStudy (anonymous):

Please answer! Find distance from -5+4i to 0

OpenStudy (anonymous):

wouldn't you just set it equal to 0? If that is the case it is 5/4

OpenStudy (dumbcow):

use distance formula d^2 = (-5)^2 + 4^2

OpenStudy (radar):

\[\sqrt{(-5^{2}+(4i)^{2}}\] would this do it?

OpenStudy (radar):

\[\sqrt{25+(-16)}=\sqrt{9}=\pm3\]

OpenStudy (anonymous):

But it asks for the distance from 0 to (-5+4i) is that still right

OpenStudy (radar):

It is a complex number containing a real component (-5) and the imaginary component 4i

OpenStudy (radar):

Helllllp

OpenStudy (anonymous):

I get it now! Thanks

myininaya (myininaya):

sqrt(5^2+4^2)=sqrt(25+16)=sqrt(41)

OpenStudy (radar):

I see where I went astray, I inludec i squared which is a -1, when I should of ignored the fact there was an i, as I was actually getting the absolute value.